rewrite 14x-7y=21, so that y is a function of x
rewrite 8x+2y=14, so that y is a function of x

Answer by Alan3354(69443)

rewrite 14x-7y=21, so that y is a function of x
rewrite 8x+2y=14, so that y is a function of x
-----------------------
The idea is to get y on one side of the equal sign by itself. Both of these eqns can be reduced, also. Most teachers expect that to be done.
---------------
14x-7y=21
2x - y = 3
y = 2x - 3
--------------
8x+2y=14
4x+y = 7
y = -4x + 7